OpenStudy (anonymous):

Simplify. tan^2(x)-sin^2(x)

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \tan^2(x)-\sin^2(x) }\) \(\LARGE\color{black}{ \frac{\sin^2(x)}{\cos^2{x}}-\frac{\sin^2(x)}{1} }\) can you do it from here?

OpenStudy (solomonzelman):

you believe it is not going to be simpler once you add the fractions? Just do what I say step by step. 1)Add the fractions after finding the denominator. (for now)

OpenStudy (solomonzelman):

(multiply top and bottom of the second fraction times \(\large\color{black}{ \cos^2{x} }\) .

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{\sin^2(x)}{\cos^2(x)}-\frac{\sin^2(x)~~ \color{red}{\times\cos^2(x) } }{1 \color{red}{~~\times \cos^2(x)}} }\) \(\LARGE\color{black}{ \frac{\sin^2(x)}{\cos^2(x)}-\frac{\sin^2(x)~~ {\cos^2(x) } }{ \cos^2(x)} }\)

OpenStudy (solomonzelman):

Can you add the fractions?